Scott is an accountant who purchased a vacant block of land in

Brisbane on 1 October 1980. On 1 September 1986, Scott built a

house on the land. At the time, the land was valued at $90,000 and

the cost of construction was $60,000. The property has been rented

out since construction was completed. On 1 March of the current tax

year, Scott sold the property at auction for $800,000.

Requirement:

a) Based on the information above, determine Scott’s net

capital gain or net capital loss for the year ended

30 June of the current tax year.

b) How would your answer to (a) differ if Scott sold the

property to his daughter for $200,000?

c) How would your answer to (a) differ if the owner of the

property was a company instead of an individual?
1
Expert's answer
2016-04-26T08:39:06-0400
The land was valued at $90,000 and the cost of construction was $60,000. The property has been rented out for 30 years. On 1 March of the current tax year, Scott sold the property at auction for $800,000.
Requirement:
a) Based on the information above, Scott’s net capital gain for the year ended 30 June of the current tax year is NCG = 800,000 - 90,000 - 60,000 = $650,000.
b) if Scott sold the property to his daughter for $200,000, then NCG = 200,000 - 90,000 - 60,000 = $50,000.
c) if the owner of the property was a company instead of an individual, then amortization payments and taxes should be deducted.